OpenStudy (anonymous):

APPLICATION OF QUADRATIC EQUATIONS The manager of a bicycle shop has found that at a price (in dollars ) of p(x)= 150 - x/5 per bicycle, x bicycles will be sold. a) Find an expression for the total revenue from the sale of x bicycles(hint: Revenue = demand x price) b)Find the number of bicycles sales that leads to maximum revenue c)find the maximum revenue. @timo86m

OpenStudy (anonymous):

hmm i suppose the demand would be represented by x so 150x- x^2/5 as far as max and min take the derivative so 150-2/5x solve for x then plug back in for revenue! so b=375 c= 56250-28125= 28125

OpenStudy (anonymous):

thanks but your solution is not really clear.if you can elaborate will be soo glad.thanks

OpenStudy (anonymous):

where you got the 150-2/5x is where im confused

OpenStudy (anonymous):

thats the derviative of 150x-(x^2)/5

OpenStudy (anonymous):

so how do i go about solving for b

OpenStudy (anonymous):

okay so by now you probably no that the derivative represents the slope correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so when the slope =0 we assume its a max or min

OpenStudy (anonymous):

since there is only one solution when i set the derivative to zero I assumed it was the max you can text for it through.

OpenStudy (anonymous):

test*

OpenStudy (anonymous):

so that x that we found is the number of bicycles being sold at that max

OpenStudy (anonymous):

which was 375

OpenStudy (anonymous):

I can explain this much easier on skype if you want

OpenStudy (anonymous):

aww ma skype aint working now .the issue is the 375 NOW HOW YOU GOT THAT

OpenStudy (anonymous):

okay so set the derivative to 0 and solve for x

OpenStudy (anonymous):

BECAUSE OF THE X SQUARED I GET 750X - X^2=0

OpenStudy (anonymous):

that is where i get to and then get confused

OpenStudy (anonymous):

i have no idea how you end up with 750x-x^2

OpenStudy (anonymous):

the derivative of 150x-(x^2)/5 is 150-2x/5

OpenStudy (anonymous):

the first equation i just listed is the revenue

OpenStudy (anonymous):

so how did you get 150-2x/5 how did u work to get that

OpenStudy (anonymous):

n(x)(n-1)

OpenStudy (anonymous):

n(x)^(n-1)*

OpenStudy (anonymous):

so 150*1(x)^(1-1) n=1 in this one and 1/5*2x^(2-1) n=2

OpenStudy (anonymous):

ok im tryna figure it out.will contact you if i have any question,.thanks.you soo helpful

OpenStudy (anonymous):

let me know

OpenStudy (anonymous):

the equation I listed needs to be put in vertex form

OpenStudy (anonymous):

sorry not vertex form im dumb. standard form. which it is. -x^2+750x the axis of symmetry is what you are looking for

OpenStudy (anonymous):

ax^2+bx+c so -b/2a, -750/2(-1)= 375